A company uses the fomula C(x)= 0.02x^2-3.4x+150 to model the unit cost in dollars for producing x stabilizer bars. For what number of bars is the unit cost at its maximum? What is the unit of cost at the level of production?
Answer by stanbon(75887)

A company uses the fomula C(x)= 0.02x^2-3.4x+150 to model the unit cost in dollars for producing x stabilizer bars. For what number of bars is the unit cost at its maximum? What is the unit of cost at the level of production?
----------
maximum occurs when x = -b/2a = 3.4/(2*0.02) = 85
Unit cost is then C(85)= 5.5
==================
Cheers,
Stan H.